In the lab, Lashonda has two solutions that contain alcohol and is mixing them with each other. She uses 4 times as much Solution A as Solution B. Solution A is 13% alcohol and Solution B is 18% alcohol. How many milliliters of Solution B does she use, if the resulting mixture has 350 milliliters of pure alcohol?

Found 2 solutions by stanbon, ankor@dixie-net.com:
Answer by stanbon(75887)

In the lab, Lashonda has two solutions that contain alcohol and is mixing them with each other. She uses 4 times as much Solution A as Solution B. Solution A is 13% alcohol and Solution B is 18% alcohol. How many milliliters of Solution B does she use, if the resulting mixture has 350 milliliters of pure alcohol?
-----
A = 4B or B = A/4
----
0.13A + 0.18B = 350ml
---------------------------
Substitute for "A" and solve for "B"::
0.13*4B + 0.18B = 350
0.52B + 0.18B = 350
0.7B = 350
B = 35000/7
B = 5000 ml = 5 litres (amt. of solution B needed)
-----------------
Cheers,
Stan H.
-------------

Answer by ankor@dixie-net.com(22740)

Lashonda has two solutions that contain alcohol and is mixing them with each other.
She uses 4 times as much Solution A as Solution B.
Solution A is 13% alcohol and Solution B is 18% alcohol.
How many milliliters of Solution B does she use, if the resulting mixture has 350 milliliters of pure alcohol?
:
let b = amt of solution B required
"She uses 4 times as much Solution A as Solution B." therefore:
a = 4b
:
Amt of alcohol in A + amt of alcohol in B = 350 ml of alcohol
.13(4b) + .18b = 350
.52b + .18b = 350
.7b = 350
b = 350/.7
b = 500 ml of B solution required